# Bayesian Controlled FDR Variable Selection via Knockoffs

### v2

@@ Abstract
In many research fields, researchers aim to identify significant associations between a set of explanatory variables and a response while controlling the false discovery rate (FDR). The Knockoff filter has been recently proposed in the frequentist paradigm to introduce controlled noise in a model by cleverly constructing copies of the predictors as auxiliary variables. In this paper, we develop a fully Bayesian generalization of the classical model-X knockoff filter for normally distributed covariates. In our approach we consider a joint model of the covariates and the response variables, and incorporate the conditional independence structure of the covariates into the prior distribution of the auxiliary knockoff variables. We further incorporate the estimation of a graphical model among the covariates, leading to improved knockoffs generation and estimation of the covariate effects on the response. We use a modified spike-and-slab prior on the regression coefficients, which avoids the increase of the model dimension as typical in the classical knockoff filter. Our model performs variable selection using an upper bound on the posterior probability of non-inclusion. We show how our construction leads to valid model-X knockoffs and demonstrate that the proposed variable selection procedure leads to controlling the Bayesian FDR at an arbitrary level, in finite samples, if the distribution of the covariates is fully known, and asymptotically if estimated as in the proposed model. We use simulated data to demonstrate that our proposal increases the stability of the selection with respect to classical knockoff methods, as it relies on the entire posterior distribution of the knockoff variables instead of a single sample. With respect to Bayesian variable selection methods, we show that our selection procedure achieves comparable or better performances, while maintaining control over the FDR. Finally, we show the usefulness of the proposed model with an application to real data.
Keywords: False Discovery Rate, Graphical Models, MCMC, Regression, Spike-and-Slab Priors.
@@ 1 Introduction
In recent years, variable selection has emerged as a crucial task at the forefront of statistics and machine learning, since it enables searching for associations between an outcome variable and a large amount of possible explanatory variables, when only a few of them are truly relevant. Hence, there is a need for methods that can accurately identify the subset of variables that affect the outcome. Classical approaches for variable selection typically rely on constrained optimization of the likelihood function (see, among others, Tibshirani,, 1996; Fan and Li,, 2001; Zou and Hastie,, 2005). In the Bayesian approach, variable selection can be performed by employing several classes of prior distributions (Tadesse and Vannucci,, 2021, for a review, see). For regression models, spike-and-slab priors, proposed for the first time by Mitchell and Beauchamp, (1988), impose mixture distributions on the regression coefficients, with latent binary indicators that identify the relevant (active) predictors (George and McCulloch,, 1997; Brown et al.,, 1998). Another approach involves the use of shrinkage priors to adaptively shrink coefficients towards zero (Park and Casella,, 2008; Carvalho et al.,, 2010).
@@
A fundamental challenge in variable selection is controlling the False Discovery Rate (FDR), to ensure the reliability and replicability of research findings. False discoveries refer to situations where a statistical procedure incorrectly identifies a variable as relevant when, in reality, its effect is merely due to random chance or noise in the data. Several multiple testing procedures have been proposed to control the FDR - that is, the expected proportion of inactive variables mistakenly identified as active, among all those selected as active (see, for instance, Benjamini and Hochberg,, 1995; Benjamini and Yekutieli,, 2001, 2005). Those procedures typically rely on computing $p$-values, which is not always an easy task (Candès et al.,, 2018). Alternative methodologies have been developed by Xing et al., (2023), who propose a method that leverages a pair of mirror variables for each predictor to provide asymptotic control on the FDR, and by Dai et al., (2022), who employ a data-splitting procedure to accomplish the same goal with minimal computational effort.
@@
In Bayesian settings, the usual definition of FDR cannot be used, since both the set of active and inactive variables are considered random. Instead, the Bayesian FDR (BFDR) is defined as an expected FDR conditionally upon the observed data, i.e., computed with respect to the posterior distribution of the sets of active/inactive variables given the observed data (Muller et al.,, 2006; Whittemore,, 2007). An advantage of Bayesian methods is that they can facilitate multiplicity adjustments through the choice of the prior. For example, in the context of variable selection, Scott and Berger, (2010) show that a spike-and-slab prior with a Beta-Bernoulli prior on the selection indicator achieves an asymptotic multiplicity control. Muller et al., (2006) and Guindani et al., (2009) prove control of the Bayesian FDR by characterizing the FDR problem as a decision problem where the optimal decision rule is based on thresholding the marginal posterior probabilities. More recently, Castillo and Roquain, (2020) prove control of the Bayes FDR for general empirical Bayes multiple testing procedures. In practical settings, the optimal threshold for controlling the FDR corresponds to the minimum marginal probability of inclusion that ensures that the BFDR is less than a desired level (Newton et al.,, 2004). Therefore, this procedure relies on the estimates of these probabilities, which are typically obtained from the MCMC output.
@@
Recent classical techniques to control the FDR do so by introducing carefully designed perturbations to the data. For example, Barber and Candès, (2015) propose the knockoff filter, a powerful method that achieves finite sample FDR control in fixed-design, low-dimensional linear models. The key idea is to add controlled noise in the form of cleverly constructed copies of the predictors. These copies, called knockoffs, are constructed to mimic the correlation structure of the true covariates while being statistically independent of the response. A linear model is then fitted to this augmented design, and variable selection is performed using a function of the coefficients of both the original and the knockoff variables. Essentially, knockoff variables act as negative controls. This framework has been extended to high-dimensional fixed-design settings in Barber and Candès, (2019), and is often referred to as the fixed-X knockoff procedure. The model-X knockoff filter of Candès et al., (2018) further extends the approach, by avoiding assumptions on the relation between the response and the covariates. Instead, it requires full knowledge of the joint distribution of the covariates. Alternative perspectives have been investigated by Sesia et al., (2018), who provide an efficient exact sampling of the knockoff variables when the covariates are distributed as a hidden Markov model, and by Bates et al., (2021), who apply a modified version of the algorithm proposed by Candès et al., (2018) to sample knockoffs exactly in the case of known independence structure of the covariates. More recently, Berti et al., (2023) provide a new characterization of the joint distribution of the observed covariates and the knockoffs, which they define in terms of copulas. When the covariates follow a joint Gaussian distribution, their approach recovers the same solution as in Candès et al., (2018).
@@
One of the main limitations of the classical model-X knockoff filter is that it depends on the quality of a single random draw of knockoff variables: different draws, from the same set of predictors, may yield different selected sets. To reduce this variability, Ren et al., (2023) combine the knockoff filter with the stability-selection framework of Meinshausen and Bühlmann, (2010), yielding a more stable selection of active variables.
In this work, we take a fully Bayesian approach that generalizes the model-X knockoff filter and builds stability directly into the inference process. More specifically, we treat the knockoffs as latent variables and update them as part of the posterior inference. We consider the joint model of the covariates and the response, and incorporate the conditional independence structure of the covariates into the prior distribution of the knockoff variables. Our joint modeling construction employs a likelihood decomposition via auxiliary variables that allows us to obtain valid model-X knockoffs.
We further assume a Gaussian graphical model to model the covariates’ dependence structure and impose a continuous spike-and-slab prior on the precision matrix (Wang,, 2015). Estimating a graphical model on the covariates allows us to obtain a coherent model that does not require perfect knowledge of the precision matrix. Moreover, unlike plug-in estimators, our Bayesian approach can naturally incorporate prior information about the conditional dependence structure of the covariates into the model. As argued by Barber et al., (2020), this leads to higher-quality knockoffs.
@@
For the regression coefficients, we adopt the mixture spike-and-slab prior suggested by Candès et al., (2018), which avoids the increase of the model dimension typical of the classical knockoff filter. Additionally, we modify this prior by allowing the probability of inclusion of each variable to depend on the number of its neighbors in the graph. We ensure that the knockoff variables possess a priori the required properties outlined in Candès et al., (2018) and demonstrate that this characterization is sufficient for controlling the BFDR at an arbitrary level, in finite samples, provided the distribution of the covariates is fully known. Moreover, for a sufficiently large sample size, our procedure maintains asymptotic FDR control when the true covariates’ graph is estimated from data, as in our proposed framework. In applications, to maintain control over the BFDR, we employ a greedy algorithm, similar to the one presented in Newton et al., (2004), but relying on an estimate of the upper bound for the probability of non-inclusion. Using simulated data, we show that our proposal substantially improves the stability of the selection compared to classical knockoff methods, as it leverages the entire posterior distribution of the knockoff variables rather than relying on a single realization. Furthermore, we show that our procedure achieves performance comparable to or better than standard Bayesian variable selection methods, while consistently controlling the FDR.
@@
To our knowledge, only a few Bayesian knockoff filters have been proposed in the literature to date. Gu and Yin, (2021) first considered a generalized linear model framework and assumed the distribution of the covariates to be completely known. With respect to their model formulation, in our approach, the incorporation of the estimation of a graphical model on the covariates aids in knockoff generation and improves the estimation of the covariate effects on the response. Furthermore, unlike the model formulation of Gu and Yin, (2021), our joint model ensures valid model-X knockoffs. Yap and Gauran, (2023) employ the Bayesian lasso of Park and Casella, (2008) for the estimation of the regression coefficients. However, they treat the covariates as fixed and require the augmented knockoff matrix to be generated prior to Bayesian inference.
@@ 2.1 A review of the Knockoff filter
Knockoff filter is a framework that enables variable selection in an FDR-controlled environment. The main idea, originally proposed by Barber and Candès, (2015), is to perturb the model by adding a set of inactive predictors built in such a way that they resemble the dependence structure of the original covariates. In Barber and Candès, (2015) the knockoff predictors were built deterministically, while Candès et al., (2018) extended the framework to the model-$X$ knockoff filter, with random covariates. The model-$X$ knockoff assumes a joint distribution $p(X)$ for the predictors and random knockoff variables, $\widetilde{X}=(\widetilde{X}_{1},\dots,\widetilde{X}_{p})$, which are defined in a way to make it hard for the model to distinguish them from $X$. Formally, the knockoffs are sampled from a distribution $p(\widetilde{X}\mid X)$ such that these two properties hold:
@@
where $(X,\widetilde{X})_{\text{Swap}(j)}$ results from swapping the $j$-th variable with the $j$-th knockoff. Note that while the first property is easily obtained, the second presents some difficulties (Sesia et al.,, 2018; Bates et al.,, 2021; Berti et al.,, 2023). A general algorithm to sample knockoffs is given in Candès et al., (2018), but it requires the computation of the conditional probability distribution of each covariate given all the others and the previously sampled knockoffs, $p(X_{j}|X_{\setminus j},\widetilde{X}_{1:j-1})$, which can be difficult to obtain for general distributions. On the other hand, when all covariates are pairwise independent, each knockoff can be sampled from the marginal distribution of $X_{j}$.
In the classical knockoff framework as proposed by Barber and Candès, (2015), given the sampled knockoff variables $\widetilde{\bm{X}}=(\widetilde{x}_{1},\dots,\widetilde{x}_{n})$, any model which returns a variable importance metric is fit using the augmented set of predictors $[\bm{X},\widetilde{\bm{X}}]$. Variable selection is performed by defining a feature statistic $W_{j}$, $j=1,\ldots,p$, which combines the variable importance metric of each original predictor with the one of its own knockoff, using any antisymmetric function such that it has a symmetric distribution around zero for inactive variables. The final set of selected variables consists of those exceeding a pre-defined threshold, chosen according to a desired FDR level.
The procedure of model-X knockoffs (Candès et al.,, 2018) described above can be simplified if the distribution of the covariates is Gaussian. Let $X\sim N(0,\bm{\Sigma})$, with $\bm{\Sigma}$ the covariance matrix, then a joint distribution for which pairwise exchangeability holds is
@@
Note that Equation (3) implies
$\widetilde{X}=X(I-\bm{\Sigma}^{-1}\text{diag}(s))+U.$ (4)
with $U\sim N(0,\bm{A})$.
@@ 2.2 Constructing Bayesian knockoffs
We now introduce our Bayesian generalization of the model-X knockoff filter. We are interested in the joint model for $Y,X$, e.g. $p(Y,X)$. In particular, our interest focuses on the elements of the factorization of $p(Y,X)$ into the conditional distribution $p(Y|X)$ and the marginal distribution $p(X)$. For the latter we assume a multivariate Gaussian model as
$X\sim N(0,\bm{\Sigma}),$ (5)
which implies that the joint distribution of $(X,\widetilde{X})$ is as reported in Equation (1). We then derive the conditional distribution $p(Y\mid X)$ as the marginal distribution over the knockoff variables $\widetilde{X}$ of the distribution $p(Y\mid X,\widetilde{X})$ as

### v3

# Bayesian Controlled FDR Variable Selection via Parameter-Expanded Latent Knockoffs

@@ Abstract
In many research fields, researchers aim to identify significant associations between a set of explanatory variables and a response while controlling the false discovery rate (FDR). The Knockoff filter has been recently proposed in the frequentist paradigm to introduce controlled noise in a model by cleverly constructing copies of the predictors as auxiliary variables. In this paper, we develop a fully Bayesian generalization of the classical model-X knockoff filter for normally distributed covariates. In our approach, we consider a joint model for the covariates and the response, where the conditional independence structure of the covariates is captured through a Gaussian graphical model and used to define a latent knockoff layer through a parameter-expanded representation of the response model. Estimating the covariate graph informs the knockoff construction and improves inference on the covariate effects. We use a modified spike-and-slab prior on the regression coefficients, which avoids the increase of the model dimension typical of the classical knockoff filter. We also address extensions to settings with non-Gaussian responses. Our model performs variable selection using an upper bound on the posterior probability of non-inclusion. We show that the induced latent knockoff layer defines valid Gaussian model-X knockoffs under the proposed construction and that the resulting procedure controls the Bayesian FDR at an arbitrary level, in finite samples, if the distribution of the covariates is fully known; under an estimated graphical structure, it satisfies an asymptotic FDR guarantee. We use simulated data to demonstrate that our proposal increases the stability of the selection with respect to classical knockoff methods, as it relies on the entire posterior distribution of the latent knockoff variables instead of a single sample. With respect to Bayesian variable selection methods, we show that our selection procedure achieves comparable or better performances, while maintaining control over the FDR. Finally, we show the usefulness of the proposed model with an application to real data.
Keywords: False Discovery Rate, Graphical Models, MCMC, Regression, Spike-and-Slab Priors
@@ 1 Introduction
In recent years, variable selection has emerged as a crucial task at the forefront of statistics and machine learning, since it enables searching for associations between an outcome variable and a large amount of possible explanatory variables, when only a few of them are truly relevant. Hence, there is a need for methods that can accurately identify the subset of variables that affect the outcome. Classical approaches for variable selection typically rely on constrained optimization of the likelihood function (see, among others, Tibshirani, 1996; Fan and Li, 2001; Zou and Hastie, 2005). In the Bayesian approach, variable selection can be performed by employing several classes of prior distributions (Tadesse and Vannucci, 2021, for a review, see). For regression models, spike-and-slab priors, proposed for the first time by Mitchell and Beauchamp (1988), impose mixture distributions on the regression coefficients, with latent binary indicators that identify the relevant (active) predictors (George and McCulloch, 1997; Brown et al., 1998). Another approach involves the use of shrinkage priors to adaptively shrink coefficients towards zero (Park and Casella, 2008; Carvalho et al., 2010).
@@
A fundamental challenge in variable selection is controlling the False Discovery Rate (FDR) to ensure the reliability and replicability of research findings. False discoveries refer to situations where a statistical procedure incorrectly identifies a variable as relevant when, in reality, its effect is merely due to random chance or noise in the data. Several multiple testing procedures have been proposed to control the FDR - that is, the expected proportion of inactive variables mistakenly identified as active, among all those selected as active (see, for instance, Benjamini and Hochberg, 1995; Benjamini and Yekutieli, 2001, 2005). Those procedures typically rely on computing $p$-values, which is not always an easy task (Candès et al., 2018). Alternative methodologies have been developed by Xing et al. (2023), who propose a method that leverages a pair of mirror variables for each predictor to provide asymptotic control on the FDR, and by Dai et al. (2022), who employ a data-splitting procedure to accomplish the same goal with minimal computational effort.
@@
In Bayesian settings, the usual definition of FDR cannot be used, since both the set of active and inactive variables are considered random. Instead, the Bayesian FDR (BFDR) is defined as an expected FDR conditionally upon the observed data, i.e., computed with respect to the posterior distribution of the sets of active/inactive variables given the observed data (Muller et al., 2006; Whittemore, 2007). An advantage of Bayesian methods is that they can facilitate multiplicity adjustments through the choice of the prior. For example, in the context of variable selection, Scott and Berger (2010) show that a spike-and-slab prior with a Beta-Bernoulli prior on the selection indicator achieves an asymptotic multiplicity control. Muller et al. (2006) and Guindani et al. (2009) prove control of the Bayesian FDR by characterizing the FDR problem as a decision problem where the optimal decision rule is based on thresholding the marginal posterior probabilities. More recently, Castillo and Roquain (2020) prove control of the Bayes FDR for general empirical Bayes multiple testing procedures. In practical settings, the optimal threshold for controlling the FDR corresponds to the minimum marginal probability of inclusion that ensures that the BFDR is less than a desired level (Newton et al., 2004). Therefore, this procedure relies on the estimates of these probabilities, which are typically obtained from the MCMC output.
@@
Recent classical techniques to control the FDR do so by introducing carefully designed perturbations to the data. For example, Barber and Candès (2015) propose the knockoff filter, a powerful method that achieves finite sample FDR control in fixed-design, low-dimensional linear models. The key idea is to add controlled noise in the form of cleverly constructed copies of the predictors. These copies, called knockoffs, are constructed to mimic the correlation structure of the true covariates while being statistically independent of the response. A linear model is then fitted to this augmented design, and variable selection is performed using a function of the coefficients of both the original and the knockoff variables. Essentially, knockoff variables act as negative controls. This framework has been extended to high-dimensional fixed-design settings in Barber and Candès (2019), and is often referred to as the fixed-X knockoff procedure. The model-X knockoff filter of Candès et al. (2018) further extends the approach, by avoiding assumptions on the relation between the response and the covariates. Instead, it requires full knowledge of the joint distribution of the covariates. Alternative perspectives have been investigated by Sesia et al. (2018), who provide an efficient exact sampling of the knockoff variables when the covariates are distributed as a hidden Markov model, and by Bates et al. (2021), who apply a modified version of the algorithm proposed by Candès et al. (2018) to sample knockoffs exactly in the case of known independence structure of the covariates. More recently, Berti et al. (2023) provide a new characterization of the joint distribution of the observed covariates and the knockoffs, which they define in terms of copulas. When the covariates follow a joint Gaussian distribution, their approach recovers the same solution as in Candès et al. (2018).
@@
One of the main limitations of the classical model-X knockoff filter is that it depends on the quality of a single random draw of knockoff variables: different draws, from the same set of predictors, may yield different selected sets. To reduce this variability, Ren et al. (2023) combine the knockoff filter with the stability-selection framework of Meinshausen and Bühlmann (2010), yielding a more stable selection of active variables.
In this work, we take a fully Bayesian approach that generalizes the model-X knockoff filter and builds stability directly into the inference process. Rather than conditioning on a single realized knockoff matrix, we introduce a latent knockoff layer through a parameter-expanded representation of a baseline response model and propagate uncertainty in that layer through posterior inference. We further assume a Gaussian graphical model to model the covariates’ dependence structure and impose a continuous spike-and-slab prior on the precision matrix (Wang, 2015). Estimating a graphical model on the covariates allows us to obtain a coherent model that does not require perfect knowledge of the precision matrix. Moreover, unlike plug-in estimators, our Bayesian approach can naturally incorporate prior information about the conditional dependence structure of the covariates into the model. As argued by Barber et al. (2020), this leads to higher-quality knockoffs. For the regression coefficients, we adopt the mixture spike-and-slab prior suggested by Candès et al. (2018), which avoids the increase of the model dimension typical of the classical knockoff filter. Additionally, we modify this prior by allowing the probability of inclusion of each variable to depend on the number of its neighbors in the graph. We also discuss extensions of the proposed Bayesian knockoff filter formulation to settings with non-Gaussian responses, focusing in particular on survival outcomes through accelerated failure time models.
@@
We ensure that the latent knockoff variables possess a priori the required properties outlined in Candès et al. (2018) and demonstrate that this characterization is sufficient for controlling the BFDR at an arbitrary level, in finite samples, provided the distribution of the covariates is fully known. When the covariate graph is estimated from data, the graphical prior yields accurate posterior estimation of the precision matrix, and the resulting procedure retains an asymptotic FDR guarantee. In applications, to maintain control over the BFDR, we employ a greedy algorithm, similar to the one presented in Newton et al. (2004), but relying on an estimate of the upper bound for the probability of non-inclusion. Using simulated data, we show that our proposal substantially improves the stability of the selection compared to classical knockoff methods, as it leverages the entire posterior distribution of the knockoff variables rather than relying on a single realization. Furthermore, we show that our procedure achieves performance comparable to or better than standard Bayesian variable selection methods, while consistently controlling the FDR.
@@
To our knowledge, only a few Bayesian knockoff filters have been proposed in the literature to date. Gu and Yin (2021) first considered a generalized linear model framework and assumed the distribution of the covariates to be completely known. In contrast, our formulation combines estimation of a Gaussian graphical model for the covariates with a parameter-expanded latent knockoff construction, so that the estimated covariate graph informs knockoff construction and inference on the regression effects. Furthermore, unlike the formulation of Gu and Yin (2021), our construction defines latent knockoff variables that constitute valid model-X knockoffs a priori. Yap and Gauran (2023) employ the Bayesian lasso of Park and Casella (2008) for the estimation of the regression coefficients. However, they treat the covariates as fixed and require the augmented knockoff matrix to be generated prior to Bayesian inference.
@@ 2.1 A review of the Knockoff filter
Knockoff filter is a framework that enables variable selection in an FDR-controlled environment. The main idea, originally proposed by Barber and Candès (2015), is to perturb the model by adding a set of inactive predictors built in such a way that they resemble the dependence structure of the original covariates. In Barber and Candès (2015) the knockoff predictors were built deterministically, while Candès et al. (2018) extended the framework to the model-$X$ knockoff filter, with random covariates. The model-$X$ knockoff assumes a joint distribution $p(X)$ for the predictors and random knockoff variables, $\widetilde{X}=(\widetilde{X}_{1},\dots,\widetilde{X}_{p})$, which are defined in a way to make it hard for the model to distinguish them from $X$. Formally, the knockoffs are sampled from a distribution $p(\widetilde{X}\mid X)$ such that these two properties hold:
@@
where $(X,\widetilde{X})_{\text{Swap}(j)}$ results from swapping the $j$-th variable with the $j$-th knockoff. Note that while the first property is easily obtained, the second presents some difficulties (Sesia et al., 2018; Bates et al., 2021; Berti et al., 2023). A general algorithm to sample knockoffs is given in Candès et al. (2018), but it requires the computation of the conditional probability distribution of each covariate given all the others and the previously sampled knockoffs, $p(X_{j}|X_{\setminus j},\widetilde{X}_{1:j-1})$, which can be difficult to obtain for general distributions. On the other hand, when all covariates are pairwise independent, each knockoff can be sampled from the marginal distribution of $X_{j}$.
The procedure of model-X knockoffs (Candès et al., 2018) described above can be simplified if the distribution of the covariates is Gaussian. Let $X\sim N(0,\bm{\Sigma})$, with $\bm{\Sigma}$ the covariance matrix, then a joint distribution for which pairwise exchangeability holds is
@@
In the classical knockoff framework as proposed by Barber and Candès (2015), given the sampled knockoff variables $\widetilde{\bm{X}}=(\widetilde{x}_{1},\dots,\widetilde{x}_{n})$, any model which returns a variable importance metric is fit using the augmented set of predictors $\bm{X},\widetilde{\bm{X}}$]. Variable selection is performed by defining a feature statistic $W_{j}$, $j=1,\ldots,p$, which combines the variable importance metric of each original predictor with the one of its own knockoff, using any antisymmetric function such that it has a symmetric distribution around zero for inactive variables. The final set of selected variables consists of those whose $W_{j}$ exceeds a pre-defined threshold, chosen according to a desired FDR level.
@@ 2.2 Constructing Bayesian knockoffs
We now introduce our Bayesian generalization of the model-X knockoff filter. We are interested in the joint model for $Y,X$, i.e., $p(Y,X)$. In particular, our interest focuses on the elements of the factorization of $p(Y,X)$ into the conditional distribution of the response $p(Y|X)$ and the marginal distribution of the covariates, $p(X)$. For the latter we assume a multivariate Gaussian model as
$X\sim N(0,\bm{\Sigma}).$ (4)
Unlike the conventional model-X knockoff procedure, where a knockoff matrix is generated once and then treated as a fixed augmentation of the design matrix, in our formulation the knockoff variables are introduced as latent auxiliary variables in a parameter-expanded working model used to construct feature scores. This representation induces a paired competition between each original covariate and its matched knockoff and allows uncertainty in the knockoff layer to be propagated through posterior inference. More precisely, we take as baseline observed-data model for the response the Gaussian regression model
$Y\mid X,\alpha,\tau^{2}\sim N\left(X\alpha,\tau^{2}I_{n}\right),$ (5)
where $\alpha$ denotes the effective regression coefficient and $\tau^{2}$ the marginal error variance. To induce the competition between each original covariate and its auxiliary knockoff, we introduce a parameter expansion, such that the effective parameters $\alpha$ and $\tau^{2}$ are reparametrized as
$\alpha=\beta+\left(I-\bm{\Sigma}^{-1}\text{diag}(s)\right)\tilde{\beta},\quad\text{and}\quad\tau^{2}=\sigma^{2}+\tilde{\beta}^{T}A\tilde{\beta}$, (6)
where $\beta$ and $\tilde{\beta}$ are $p$-dimensional vectors. Parameter expansion has often been used to augment a model with auxiliary parameters and latent variables while leaving the observed-data model unchanged (Liu et al., 1998; Gelman, 2004). Here, we follow the same principle to obtain a reparametrization that preserves the observed-data model and underlies the construction of an antisymmetric feature statistic for variable selection. This reparametrization is motivated by the fact that it admits a convenient complete-data representation. Specifically, let $U$ be an $n\times p$ latent matrix with independent rows $u_{i}\sim N(0,\bm{A})$, where $\bm{A}$ is defined as in Equation (2), and define
$\widetilde{X}=X(I_{p}-\bm{\Sigma}^{-1}\text{diag}(s))+U.$ (7)
Then, conditional on $X$, the latent matrix $\widetilde{X}$ follows the Gaussian model-X knockoff distribution in Equation (3). Equivalently, the pair (X,X~)($\widetilde{X}$) follows the joint Gaussian distribution in Equation (1), and therefore preserves the usual pairwise exchangeability property. Under the reparametrization above, the baseline model in Equation (5) admits the complete-data representation
$Y=X\beta+\widetilde{X}\widetilde{\beta}+\epsilon=X\left[\beta+\left(I_{p}-\bm{\Omega}\operatorname{diag}(s)\right)\widetilde{\beta}\right]+U\widetilde{\beta}+\epsilon,$ (8)
where $\epsilon\sim N(0,\sigma^{2}I_{n})$ and $\boldsymbol{\Omega}=\boldsymbol{\Sigma}^{-1}$ is the precision matrix. Integrating out $U$ recovers the observed-data model with effective parameters $\alpha$ and $\tau^{2}$,
$Y\mid X,\beta,\widetilde{\beta},\Omega,\sigma^{2}\sim N\!\left(X\left[\beta+\left(I_{p}-\Omega\operatorname{diag}(s)\right)\widetilde{\beta}\right],\left(\sigma^{2}+\widetilde{\beta}^{T}A\widetilde{\beta}\right)I_{n}\right),$ (9)
which coincides with Equation (5) under the reparametrization above. Therefore, in this construction, $\beta$ and $\widetilde{\beta}$ should be interpreted as auxiliary components of the expanded representation of the effective coefficient $\alpha$, rather than as two separately identifiable regression effects in the observed-data model. This expanded representation induces a paired competition between each original covariate and its matched knockoff, which we summarize through an antisymmetric feature statistic for variable selection in Section 2.3.
2.2.1 Extensions to non-Gaussian responses
Our Bayesian knockoff filter formulation can be extended to models with non-Gaussian responses that admit a latent-variable representation for inference. Examples include probit models (Albert and Chib, 1993), for binary and multinomial responses, and accelerated failure times (AFT) models (Kalbfleisch and Prentice, 2002; Wei, 1992) for survival outcomes.
The underlying principle mirrors that of the Gaussian case. Specifically, one begins by specifying a baseline observed-data model for $Y\mid X$, or, when appropriate, for a latent response conditional on $X$, and then introduces the parameter-expanded knockoff representation through the auxiliary coefficients $\widetilde{\beta}$ and the latent Gaussian perturbation $U$. This construction produces a complete-data formulation in which each original covariate is systematically paired with its knockoff analogue, while preserving the original observed-data model upon marginalization over the augmentation variables.
As an illustration, consider an AFT model for survival data, with censored observations. Let $T_{i}$ represent the time-to-event for the $i$-th subject and $c_{i}$ the censoring time. A Gaussian AFT model assumes $\log(T_{i})=X_{i}\beta+\epsilon_{i}$, where the errors $\epsilon_{i}$ are i.i.d. from some error distribution. Here, we assume the errors to be i.i.d. Gaussian variables. To accommodate censored observations, we introduce a latent variable $y_{i}$ such that